You are given the information shown in the table about production relationships in Singapore and the rest of the world.

Inputs per Unit of Cloth Output

Singapore 75

Rest of the world 50


Inputs per Unit of Rice Output

Singapore 100

Rest of the world 50

You make several Ricardian assumptions: These are the only two commodities, there are constant ratios of input to output whatever the level of output of rice and cloth, and competition prevails in all markets.

i) If no international trade is allowed, what price ratio would prevail between rice and cloth within Singapore.

ii) If free international trade is opened up, what goods will Singapore export? Import?


1
Expert's answer
2021-10-20T10:02:25-0400

i

If no international trade is allowed, the price ratio between rice and cloth within Singapore is

"price \space ratio=\frac{P_{rice}}{P_{cloth}}=\frac{input\space for\space rice}{input\space for\space cloth}"

=opportunity cost of rice

Opportunity cost of rice "=\frac{75}{100}=\frac{3}{4}"


price ratio=labor requirement ratio


ii

if free international trade is allowed

"opportunity\space cost\space of \space rice\space(in\space singapore)\le\frac{p_{rice}}{P_{cloth}}\le opportunity\space cost\space of\space rice\space (rest\space of\space the\space world)"


"0.75\le\frac{P_rice}{P_{cloth}}\le1"


Since, Singapore has a comparative advantage in rice Singapore will export rice

and Singapore will import cloth from rest of the world.